Suppose that 905 smokers each received a nicotine patch, which delivers nicotine to the bloodstream at a much slower rate thaSuppose that 905 smokers each received a nicotine patch, which delivers nicotine to the bloodstream at a much slower rate than cigarettes do. Dosage was decreased to 0 over a 12 week period. Of these 905 people, 255 were still not smoking 6 months after treatment. Assume this sample is representative of all smokers.
Answer #1

a) Sample proportion = 255/905 = 0.282

b) Met; Met

c) For 95% confidence, z = 1.96

So,

Margin of error

\small E = 1.96*\sqrt{\frac{0.282*(1-0.282)}{905}}

E = 0.029

d) Interpretation: Option A is correct.